# Watch Your Steps: Observable and Modular Chains of Thought

### v1

@@ 1 Introduction
While chain of thought (CoT) prompting is powerful, standard CoT outputs can be “unfaithful” Jacovi and Goldberg (2020): i.e., CoT can lead to incorrect (but superficially plausible) explanations for biased outputs Turpin et al. (2024), and CoT explanations “may not align with …sequential causal reasoning” Bao et al. (2024).
Although unfaithful explanations do not affect CoT’s utility as a means of improving performance of prompted models, they do reduce the potential of CoT for other purposes, e.g., justifying a response to an end user. Despite numerous proposals Lanham et al. (2023); Bentham et al. (2024); Parcalabescu and Frank (2024), unfaithfulness remains difficult to detect and measure. The potential unfaithfulness of CoT explanations presents a jarring contrast with symbolical proofs, in which modular, verifiable reasoning steps are combined in well-understood ways.
We believe that a significant obstacle to progress on the faithfulness of CoT prompts is their syntactic diversity: because a CoT explanation can take nearly any form, they are difficult to analyze in any general way. The goal of this paper is to make CoT explanations easier to analyze while preserving the power, generality and flexibility of CoT. To do this, we propose a new prompting variant of CoT prompting in which few-shot CoT demonstrations are wrapped in a semi-formal syntax which (1) identifies and names steps;(2) defines the input/output behavior of steps; and (3) replaces CoT explanations of in-context examples with chains of these formalized steps on the same examples.
@@
We use Python syntax to describe steps, and call our method Program Trace Prompting (PTP). As shown in Figure 1, CoT demonstrations are replaced with documentation for a Python program, together with traces of that program’s behavior on some sample inputs. Each Python subroutine corresponds to a different kind of “step”, and in addition to traces, the prompt includes stubs (function-level comments and type signatures) for the subroutines called in the traces. An LLM is then prompted with this information as context, and asked to produce a trace for a novel input. An answer to the question is finally extracted from the LLM-generated trace.
Note that in Program Trace Prompting, no code is presented to the LLM. Similarly, there is no “tool use”; no generation of code or pseudo-code that is executed by Python or some other engine; and no pipeline of LLM calls as in agent frameworks (e.g., LangChain Topsakal and Akinci (2023) or DsPy Khattab et al. (2023)). Instead, for a new test input, the LLM simply generates a new trace, similar to the ones given in the prompt, but appropriate for the test input.
Importantly, the program suggested by the traces typically do not exist, and in the more interesting applications of Program Trace Prompting, cannot exist as pure Python. In Figure 1, for example, the subroutine sport_for requires using background knowledge about arbitrary entities and phrases. While this cannot be implemented directly in Python, a strong LLM can readily perform the underlying task (i.e., returning the sport associated with a well-known athlete, or a phrase describing some sport-related action), and hence the LLM can generate a trace. Hence the traces presented are best thought of as traces of Python-like “pseudocode” Li et al. (2023); Weir et al. (2024).
In this paper, we describe a particular implementation of Program Trace Prompting, and demonstrate that it has these properties.
@@
Since steps in our model are represented explicitly, one can characterize errors in reasoning as local to a particular step. However, a final answer may be wrong even when every local step seems to be correct. We call these non-local errors, but a more descriptive name might be “program induction error”; intuitively, non-local errors arise because the LLM has “guessed” the wrong algorithm from the few-shot program traces.
In our analysis we show that the majority of incorrect final answers are associated with local errors, especially for the algorithmically-simpler NLP tasks.
To summarize, the contributions of this paper are (1) description of a new framework for CoT prompting that allows detailed analysis of the “steps” used in a CoT explanation; (2) demonstration of the generality of this framework by testing it on 23 diverse tasks; and (3) presentation of methods to evaluate two newly-defined aspects of CoT reasoning, namely the modularity of CoT steps and the locality of CoT errors; and (4) an analysis of modularity of steps over a large set of tasks and step types.
@@ 2.2.1 Mocks based on CoT prompts
In all our experiments, we constructed program trace prompts that closely followed the existing CoT prompt for the task (as suggested by Figure 1.) The existing CoT prompts were also used as a performance baseline. Modifications to CoT prompts were made only (1) to encourage the model to closely follow the in-context demonstrations, and in particular to produce output in an easily-parsed format and (2) rarely, to fix errors in the CoT prompt. See Appendix A.1 for details.
We used a semi-automatic scheme to produce our prompts. For each task, we manually write a Python program that implements the algorithm suggested by the CoT prompt; this is easy because the program need work only for the few in-context examples appearing in the CoT prompt. For instance, the implementation of sport_for just retrieves the required values from a small dictionary containing entries for the specific entities and phrases in the CoT prompt. Following the parlance of unit testing, we call these programs mocks. [^1] Mocks are a standard way of testing interactions with a complex system that you don’t want to actually call at testing time. Functions can be nested (recursively if necessary) in a mock. When a mock is executed it automatically prints a trace; see Appendix A.2 for more information. Our traces primarily record function calls, although they can also include printed output—notably, as in Figure 1, the trace contains printed output for the “final answer” to a question.
Our placement of the traces (in the function-level documentation of the top-level function) follows a widely used convention for documenting behavior by giving input/output examples. [^2] The Python3 doctest module supports unit testing for functions that follow this convention.
@@ 2.2.2 Mock development
Our process is semi-automatic because the mocks are written by hand. For many tasks, the mock is a thin wrapper around the steps in the CoT; however, some meaningful decisions do need to be made by mock authors. For instance, in Figure 1, the function sport_for might plausibly return a list of strings, rather than a single string, and the consistent_sports function could plausibly be omitted, as it is implicit in the CoT prompt.
To avoid overfitting (when making these decisions), we used 60 random examples of each task to design the mocks, and reserved the rest (190 examples for most tasks) for a final unseen test set set. The 60 examples are further divided into “dev” and “tuning” sets with 30 examples in each. Mock implementation decisions were made based on dev data performance, but the actual values in the mocks are based on the CoT prompts, not the dev set. The tuning set was held out for the annotation experiments of Section 3.3.
@@ 2.3.1 Prompting to perform a single step
The result of single-step prompting is more variable than the result of PTP prompting on full examples. One common failure mode for single-step prompts is that the LLM returns a trace that begins with the requested step, but then continues, often “solving” steps that might plausibly follow the requested step. We call this overgeneration.
Overgeneration can be reduced by two methods. One is to add single-step traces to the stubs for the subroutine being probed. Empirically, we find that performance for single-step prompting is improved even if the only traces added are ones that are part of the top-level traces. [^4] Specifically, improved performance is obtained by just copying over the appropriate part of the top-level CoT trace into the “stub” defining a step, as discussed in Section A.4. A second method is to use more robust ways of extracting output from single-step generations, e.g., parsing the trace to identify steps, and then using, as the single-step result, the return value from the first step with correct subroutine name. More details are given in Section 4.1 and Appendix A.4.
@@ 2.4 Language models
Program Trace Prompting can be used, in principal, with any LLM. After preliminary studies, we used Anthropic’s Sonnet 3 model (2024-06-20) as the basis for developing our prompts, as it obtains a good balance between cost, reliability, and performance. Unless we state otherwise, this is the LLM model used in the experiments below.
@@ 3.1 How well do PT prompts work?
Table 2 summarizes the accuracy results on the BBH tasks. To avoid prompt-tuning that might overfit the test data, the results in the table are generally for the first PT prompt for a task that was evaluated on the test data. (For the three starred cases, problems discovered were in the PTP prompt late in experimental procedure, and the results are for the second prompt tested–see Appendix B).
The results of Table 2 strongly support our first claim, that PTP is applicable to a broad range of tasks, in the sense that it can be used on a broad range of tasks without a major impact in performance. Of the 23 diverse tasks in BBH, switching from an informal format for explanations to a more precise one led to a statistically significant loss in performance for only one task. Three tasks showed significant improvements in performance, and average performance of PT prompts is slightly improved over the CoT baselines.
In Table 3, we also show results comparing PTP and CoT prompting for several other proprietary models, and one strong closed model (Deepseek Coder V2). These experiments were run on a subset of the data, namely the dev set for six tasks chosen to represent different kinds of tasks and different levels of accuracy. For most models the average performance is similar (within 3-4 points) for both prompting methods, but PTP appears have lower relative performance on Gemini Flash and GPT4o-Mini. todo: add/discuss open source models.
As discussed above, our experimental procedure differs from most prior work, but the CoT baseline we use is strong compared to many prior papers; see Table 11 in the Appendix for comparisons.
@@ 3.2 Are generated traces syntactically well-formed?
The main goal of Program Trace Prompting is to produce structured explanations, to enable deeper analysis of CoT performance. To verify that the explanations do indeed have the intended structure, we wrote tools to automatically evaluate generated traces for the following properties.
@@
Type correctness. Syntactically correct arguments and return values were checked to check if they were type-correct, according to the type hints given in the stubs for each step.
@@ 3.3 What kinds of errors are made?
Table 6: Number of steps and trace abstraction entropy of the 23 BBH Tasks (on the test data). Lower entropy indicates a more predictable sequence of steps. Only non-zero non-local error rates are shown.
We manually annotated all of the traces from the tune set that led to an incorrect final result, and marked the first clearly incorrect step output, if there was one, and otherwise marked the trace as containing a non-local error. [^5] Only the first local error was marked, because errors often cascade in CoT reasoning chains. The results are shown in Table 5, with more detail on non-local errors in Table 6.
@@ 3.4 Why do non-local errors occur?
To our knowledge, the existence of non-local errors in CoT reasoning has not been discussed previously in the literature. There are several possible reasons for this. First, non-local errors are relatively rare: only about 2.6% of the traces had non-local errors. Second, non-local errors occurred mainly in a small number of tasks. More than 80% of the non-local errors were in three of the tasks (Dyck Languages, Geometric Shapes, or Formal Fallacies), so more than 85% of the tasks contained very very few non-local errors. Third, non-local errors are hard to detect, even by careful manual annotators. Despite this, non-local errors are not unimportant: 22% of the incorrect traces were annotated as having non-local errors.
@@
To test this, we analyzed the traces produced for each task on the test examples. For every legal trace, we simplify the trace by removing the arguments and return values, leading to a abstract trace. For example, the traces of Figure 1 all have the same abstract trace: analyze_input() sport_for() sport_for() consistent_sports(). For several tasks, only one abstract trace is ever generated on any test example, but many have more diverse kinds of abstract traces: in particular tasks that have varying-sized inputs, or inputs of varying structure, have many distinct abstract traces.
To quantify this we computed the entropy of the empirical distribution of abstract traces for each task, which we show in Table 6. There is a clear correspondence between high trace entropy and non-local errors: the three highest non-local error rates are among the six highest-entropy tasks. The correlation between error rate and trace entropy is $r=0.51$. In contrast, the correlation to the number of steps in the trace (also shown in the table) is only $r=0.04$.
To summarize, structuring CoT explanations with Program Trace Prompting not only reveals the existence of non-local errors, but also allows us to predict (to some extent) where they might occur.
These results also emphasize a little-discussed function of CoT prompting: in addition to “unlocking reasoning” Wei et al. (2022), a CoT prompt can also provide a new sort of information to the LLM, as its steps describe a particular strategy for solving a task, which the LLM can then learn. This “strategy learning” is reliable for tasks of low and medium complexity.
@@ 4.1.1 Individual steps can be evaluated in isolation
Evaluating performance on a single step is difficult, since we do not have gold labels for each step’s output. Fortunately, several individual steps in the BBH tasks are relatively easy to encode with rules—so easy, in fact, that in writing the mocks, these steps were implemented by general Python routines, rather than demonstration-specific dictionaries. The local correctness of these steps can be readily evaluated by using the corresponding Python function from the mock as an oracle. In the BBH suite, there were 16 different steps from five tasks with oracles.
@@
We evaluated these oracle-testable steps on the inputs that were actually used in the course of solving the dev set examples (a total of 480 step executions). [^6] Note that this evaluation can be expensive, since it requires running an LLM many more times than does Program Trace Prompting, as there are an average of 11-12 steps per task. Table 7 shows the results, as well as macro-averaged overall accuracy.
The overall performance is more than 90%, so the steps are “meaningful”—the model “understands” then well enough to execute them in a “step by step” manner, where each step is performed by an independent LLM call. Here we use the full method described in Section 2.3.1, including the two methods used to address overgeneration, adding micro-traces, and parsing traces to find the first relevant output. Table 7 also shows the result of ablating these, showing that they each contribute 3-4 points of performance.
The last column of the table also shows the accuracy of the same steps when executed in the context of problem-solving—i.e, the accuracy of the steps when they occur inside a PTP trace. Note that even the best single-step prompt is significantly worse than performing the corresponding step in the context of problem-solving. [^7] Overall 10 of the 16 steps have lower average accuracy when executed independently than when executed as part of PTP problem-solving, and the remainder have the same accuracy. This is significant with $p<0.01$ with a paired $t$-test.

### v2

@@ 1 Introduction
While chain of thought (CoT) prompting is powerful, standard CoT outputs can be “unfaithful” Jacovi and Goldberg (2020): i.e., CoT can lead to incorrect but superficially plausible explanations for biased outputs Turpin et al. (2024), and CoT explanations “may not align with …sequential causal reasoning” Bao et al. (2024).
Although unfaithful explanations do not affect CoT’s utility as a means of improving performance of prompted models, they do reduce the potential of CoT for other purposes, e.g., justifying a response to an end user. Despite numerous proposals Lanham et al. (2023); Bentham et al. (2024); Parcalabescu and Frank (2024), unfaithfulness remains difficult to detect and measure. The potential unfaithfulness of CoT explanations presents a jarring contrast with symbolic proofs, in which modular, verifiable reasoning steps are combined in well-understood ways.
We believe that a significant obstacle to progress on the faithfulness of CoT prompts is their syntactic diversity; because a CoT explanation can take nearly any form, they are difficult to analyze in any general way. The goal of this paper is to make CoT explanations easier to analyze while preserving the power, generality and flexibility of CoT. To do this, we propose a new variant of CoT prompting in which few-shot CoT demonstrations are wrapped in a semi-formal syntax which (1) identifies and names steps; (2) defines the input/output behavior of steps; and (3) replaces each CoT explanation in a demonstration with an equivalent chain of formalized steps.
@@
We use Python syntax to describe steps, and call our method Program Trace Prompting (PTP). As shown in Figure 1, CoT demonstrations are replaced with documentation for a Python program, together with traces of that program’s behavior on the demonstration inputs. Each different kind of “step” is associated with a Python subroutine, and in addition to traces, the prompt includes stubs (function-level comments and type signatures) for these subroutines. An LLM is then prompted with this information as context, and asked to produce a trace for a novel input. An answer to the question is finally extracted from the LLM-generated trace.
Note that in Program Trace Prompting, no code is presented to the LLM. Similarly, there is no “tool use”; no generation of code or pseudo-code to be executed by Python or some other engine; and no pipeline of LLM calls as in agent frameworks like LangChain Topsakal and Akinci (2023) or DsPy Khattab et al. (2023). Instead, for a new test input, the LLM simply generates a new trace, similar to the ones given in the prompt, but appropriate for the test input.
Importantly, the program suggested by the traces typically do not exist, and in the more interesting applications of Program Trace Prompting, cannot exist as pure Python. In Figure 1, for example, the subroutine sport_for requires using background knowledge about arbitrary entities and phrases. While this cannot be implemented directly in Python, a strong LLM can readily perform the underlying task and thus generate a trace. Hence, the traces generated by PTP are best thought of as traces of Python-like “pseudocode” Li et al. (2023); Weir et al. (2024).
In this paper, we describe a particular implementation of Program Trace Prompting, and demonstrate that it has the following properties.
@@
Since steps in our model are explicitly represented, one can characterize errors in reasoning as local to a particular step. However, a final answer may be wrong even when every local step seems to be correct. We call these non-local errors, but a more descriptive name might be “program induction errors”; intuitively, non-local errors arise because the LLM has “guessed” the wrong algorithm from the few-shot program traces. In our analysis we show that the majority of incorrect final answers are associated with local errors, especially for the algorithmically-simpler NLP tasks.
To summarize, the contributions of this paper are (1) description of a new framework for CoT prompting that allows detailed analysis of the “steps” used in a CoT explanation; (2) demonstration of the generality of this framework by testing it on 23 diverse tasks, and releasing the corresponding prompts to the community; (3) presentation of methods to evaluate two newly-defined aspects of CoT reasoning, namely the modularity of CoT steps and the locality of CoT errors; and (4) an analysis of the modularity of steps over a large set of tasks and step types.
@@ 2.2.1 Mocks based on CoT prompts
In all our experiments, we constructed program trace prompts that closely followed the existing CoT prompt for the task (as suggested by Figure 1.) The existing CoT prompts were also used as a performance baseline. Modifications to CoT prompts were made only (1) to encourage the model to closely follow the in-context demonstrations, and in particular to produce output in an easily-parsed format and (2) rarely, to fix reasoning errors in the CoT prompt. See Appendix A.1 for details.
Program trace prompts were produced semi-automatically. For each task, we manually wrote a Python program that implements an algorithm suggested by the CoT prompt; this is feasible because the program only needs to for the few examples appearing in the CoT prompt. (For instance, the implementation of sport_for just retrieves the required values from a small dictionary containing entries for the specific entities and phrases in the CoT prompt.) Following the parlance of unit testing, we call these programs mocks. [^1] Mocks are a standard way of testing interactions with a complex system that you don’t want to actually call at testing time. Functions can be nested (recursively if necessary) in a mock. When a mock is executed it automatically prints a trace; see Appendix A.2 for more information. Our traces primarily record function calls, although they can also include printed output—notably, as in Figure 1, the trace contains printed output for the “final answer” to a question.
Our placement of the traces (in the function-level documentation of the top-level function) follows a widely used convention for documenting function behavior by giving input/output examples. [^2] The Python3 doctest module supports unit testing for functions that follow this convention.
@@ 2.2.2 Mock development
Our process is semi-automatic because the mocks are written by hand. For many tasks, the mock is a thin wrapper around the steps in the CoT demonstrations; however, some meaningful decisions do need to be made by mock authors. For instance, in Figure 1, the function sport_for might plausibly return a list of strings, rather than a single string, and the consistent_sports function could plausibly be omitted, as it is implicit in the CoT prompt.
To reduce overfitting when making these decisions, we used 60 random examples of each task to help design the mocks, and reserved the rest (190 examples for most tasks) as an unseen test set set. The 60 examples are further divided into “dev” and “tuning” sets with 30 examples in each. Mock implementation decisions were made based on dev data performance, but the actual values in the mocks are based on the CoT prompts, not the dev set. The tuning set was held out in early mock design for the annotation experiments of Section 3.3.
@@ 2.3.1 Prompting to perform a single step
The result of single-step prompting is more variable than the result of PTP prompting on full examples. One common failure mode for single-step prompts is that the LLM returns a trace that begins with the requested step, but then continues, usually tracing steps that might plausibly follow the requested step. We call this overgeneration.
Overgeneration can be reduced by two methods. One is to add single-step traces to the stubs for the subroutine being probed. Empirically, we find that performance for single-step prompting is improved even if the only traces added are ones that are part of the top-level traces. [^4] Specifically, improved performance is obtained by just copying over the appropriate part of the top-level CoT trace into the “stub” defining a step, as discussed in Section A.4. A second method is to use more a robust method for extracting output from single-step generations, i.e., parsing the trace to identify steps, and then using, as the single-step result, the return value from the first step with correct subroutine name. More details are given in Section 4.1 and Appendix A.4.
@@ 2.4 Language models
In principal, Program Trace Prompting can be used with any LLM. After preliminary studies, we used Anthropic’s Sonnet 3 model (2024-06-20) as the basis for developing our prompts, as it obtains a good balance between cost, reliability, and performance. Unless we state otherwise, this is the LLM model used in the experiments below. Program Trace Prompting was also tested on other models, as described below.
@@ 3.1 How well do PT prompts work?
Table 2 summarizes the accuracy results on the BBH tasks. To avoid prompt-tuning that might overfit the test data, the results in the table are generally [^5] For the three starred cases, problems discovered were in the PTP prompt late in experimental procedure, and the results are for the second prompt tested–see Appendix B. for the first PT prompt for a task that was evaluated on the test data.
The results of Table 2 strongly support our first claim, that PTP is applicable to a broad range of tasks, in the sense that it can be used on a broad range of tasks without a major impact in performance. Of the 23 diverse tasks in BBH, switching from an informal format for explanations to a more precise one led to a statistically significant loss in performance for only two tasks. Three tasks showed significant improvements in performance, and average performance of PT prompts is slightly improved over the CoT baselines.
As discussed above, our experimental procedure differs from most prior work, but the CoT baseline we use is strong compared to many prior papers; see Table 11 in the Appendix.
In Table 3, we also show results comparing PTP and CoT prompting for several other proprietary models, one strong open-weight model (Deepseek Coder V2, a 236B model), and several smaller models. These experiments were run on a subset of the data, namely the dev set for six tasks chosen to represent different kinds of tasks and different levels of accuracy. For most strong models the average performance is similar (within 3-4 points) for both prompting methods; however, PTP underperforms CoT prompting on many smaller models.
@@ 3.2 Are generated traces syntactically well-formed?
The main goal of Program Trace Prompting is to produce structured explanations, and thus enable deeper analysis of CoT performance. To verify that the explanations do indeed have the intended structure, we wrote tools to automatically evaluate generated traces for the following properties.
@@
Type correctness. Syntactically correct arguments and return values were checked to see if they were type-correct, according to the type hints given in the stubs for each step.
@@ 3.3 What kinds of errors are made?
Table 6: Number of steps and trace abstraction entropy of the 23 BBH Tasks (on the test data). Lower trace entropy indicates a more predictable sequence of steps. Only non-zero non-local error rates are shown.
We manually annotated all of the traces from the tune set that led to an incorrect final result. We marked the first [^6] Only the first local error was marked, because errors often cascade in CoT reasoning chains. clearly incorrect step output, if there was one, and otherwise marked the trace as containing a non-local error. The results are shown in Table 5, with more detail on non-local errors in Table 6.
@@ 3.4 Why do non-local errors occur?
To our knowledge, the existence of non-local errors in CoT reasoning has not been discussed previously in the literature. There are several possible reasons for this. First, non-local errors are relatively rare: only 2.6% of the traces had non-local errors. Second, most non-local errors occurred in a small number of tasks: more than 80% of the non-local errors were in three of the tasks (Dyck Languages, Geometric Shapes, or Formal Fallacies), so more than 85% of the tasks contained very very few non-local errors. Third, non-local errors are hard to detect, even by careful manual annotators. Despite this, non-local errors are not unimportant: 22% of the incorrect traces were annotated as having non-local errors.
@@
To test this, we analyzed the traces produced for each task on the test examples. For every legal trace, we simplify the trace by removing the arguments and return values, leading to a abstract trace. For example, the traces of Figure 1 all have the same abstract trace: analyze_input sport_for sport_for consistent_sports. For some tasks, only one abstract trace is ever generated on any test example, but many tasks have diverse abstract traces, especially tasks with inputs that vary in size or structure.
To quantify the diversity of abstract traces, we computed the entropy of the empirical distribution of abstract traces for each task, which we show in Table 6. There is a clear correspondence between high trace entropy and non-local errors: the three highest non-local error rates are among the six highest-entropy tasks. The correlation between error rate and trace entropy is $r=0.51$. In contrast, the correlation to the number of steps in the trace (also shown in the table) is only $r=0.04$.
To summarize, structuring CoT explanations with Program Trace Prompting not only reveals the existence of non-local errors, but also allows us to predict where non-local errors might occur.
These results also emphasize a little-discussed function of CoT prompting: in addition to “unlocking reasoning” Wei et al. (2022), a CoT prompt can also provide a new sort of information to the LLM, as its steps describe a particular strategy for solving a task, which the LLM can then learn. This “strategy learning” seems to be very reliable for tasks of low and medium complexity.
@@ 4.1.1 Individual steps can be evaluated in isolation
Evaluating performance on a single step is difficult, since we do not have gold labels for each step’s output. Fortunately, several individual steps in the BBH tasks are relatively easy to encode with rules—so easy, in fact, that in writing the mocks, these steps were implemented by general Python routines, rather than demonstration-specific dictionaries. The local correctness of these steps can be readily evaluated by using the corresponding Python function from the mock as an oracle.
@@
In the BBH suite, there were 16 different steps from five tasks with oracles. We evaluated these oracle-testable steps on the inputs that were actually used in the course of solving the dev set examples (a total of 480 oracle-testable step executions). [^7] Note that this evaluation can be expensive, since it requires running an LLM many more times than does Program Trace Prompting, as there are an average of 11-12 steps per task. Table 7 shows the results, as well as macro-averaged overall accuracy.
The overall performance is more than 90%, so the steps are “meaningful”—i.e., the model “understands” then well enough to execute them reliably in a “step by step” manner, where each step is performed by an independent LLM call. Here we use the full method described in Section 2.3.1, including the two methods used to address overgeneration, adding micro-traces, and parsing traces to find the first relevant output. Table 7 also shows the result of ablating these, showing that they each contribute to performance.
The last column of the table also shows the accuracy of the same steps when executed in the context of problem-solving—i.e, the accuracy of the steps when they occur inside a PTP trace. Note that even the best single-step prompt is significantly worse than performing the corresponding step in the context of problem-solving, [^8] Overall 10 of the 16 steps have lower average accuracy when executed independently than when executed as part of PTP problem-solving, and the remainder have the same accuracy. This is significant with $p<0.01$ with a paired $t$-test. which we discuss further below.